# High-Throughput Phenotyping of Clinical Text Using Large Language Models

### v1

@@ V Discussion
Nonetheless, the case for applying large language models to high-throughput phenotyping is compelling [52, 69, 50, 51, 70, 25, 71, 27]. These models are fast, accurate, and ready to run ‘out of the box.’ Unlike traditional neural network models, they do not rely on an extensive corpus of manual annotations. These models should be generalizable to a variety of diseases without additional training. Current limitations in sign normalization can be addressed using techniques from augmented retrieval generation [72], by additional pre-training, or by creating small specialized models specifically for sign normalization. Large language models such as GPT-4 are expected to become the dominant method for high-throughput clinical text phenotyping.

### v2

@@ V Discussion
Nonetheless, the case for applying LLMs to high-throughput phenotyping is compelling [37, 58, 35, 36, 59, 18, 60, 22]. These models are fast, accurate, and ready to run ‘out of the box.’ Unlike traditional neural network models, they do not rely on an extensive corpus of manual annotations. These models should be generalizable to a variety of diseases without additional training. Current limitations in sign normalization can be addressed using techniques from augmented retrieval generation [61, 40], by additional pre-training, or by creating small specialized models specifically for sign normalization. Large language models such as GPT-4 are expected to become the dominant method for high-throughput clinical text phenotyping.